Ancestry Furniture Company manufactures and sells oak tables and chairs. Price and cost data for the furniture follow: (Click1 Data Table Tables Chairs $ 100 Sales Price Variable manufacturing costs Sales commission (6%) 900 $ 621 54 Print Done0 Requirements 1. Calculate the total contribution margin and the contribution margin ratio for each sales representative (roRequirement 1. Calculate the total contribution margin and the contribution margin ratio for each sales representative (roundCalculate the total contribution margin for each sales representative. Sales representative: Albert Buck Candy Net sales reveNow calculate the contribution margin ratio for each sales representative. (Round to two decimal places—the nearest hundredthNow calculate the contribution margin ratio for each sales representative. (Round to two decimal places—the = Contribution maRequirement 2. Which sales representative has the highest contribution margin ratio? Explain why. is the sales representativehad the highest level of variable costs. had the highest sales revenue for the period. had the lowest level of fixed costs. s

PLEASE HELP FAST AND ALL IN ACCOUNTING. I WILL RATE 5 STARS.
Answer #1
1 sales representative; Tables sold Chairs per table Total chairs sold
a b a*b
Albert 60 6 360
Buck 100 8 800
Candy 110 4 440
Contribution margin=Net sales revenue-Variable costs
Sales representative; Albert Buck Candy
Net sales revenue a 90000 170000 143000
(60*900)+(360*100) (100*900)+(800*100) (110*900)+(440*100)
Variable costs;
Variable manufacturing costs 47700 85300 81070
(60*621)+(360*29) (100*621+800*29) (110*621+440*29)
Sales commission at 6%
(Net sales revenue*6%) 5400 10200 8580
Total variable costs b 53100 95500 89650
Contribution margin 36900 74500 53350
Contribution margin ratio=Contribution margin/Net sales revenue
Contribution margin ratio
Albert 36900/90000 41.00%
Buck 74500/170000 43.82%
Candy 53350/143000 37.31%
2 Buck is the sales representative with the highest contribution margin ratio because this representative
Sold the most chairs which have the higher contribution margin as compared to tables
Contribution margin ratio:
Tables (900-621)/900=279/900=0.31=31%
Chairs (100-29)/100=71/100=71%